Marilyn bought a used bike for $8 more than half its original price. marilyn paid $50.00 for the bike. What was the origianl price of the bike?
Found 2 solutions by oscargut, stanbon:
Answer by oscargut(891)

Let x the original price
then
50=8+(x/2)
then
(x/2)=42
x=84
Answer: the original price is $84

Answer by stanbon(57282)

Marilyn bought a used bike for $8 more than half its original price. marilyn paid $50.00 for the bike. What was the originall price of the bike?
--------------------------------
Let the original price be "x".
EQUATION:
(1/2)x + 8 = 50
(1/2)x = 42
x = $84 (original price)
=============================
Cheers,
Stan H.